OpenStudy (anonymous):

1.The distance traveled by an object can be modeled by the equation d = ut + 0.5at2 where d = distance, u = initial velocity, t = time, and a = acceleration. Solve this formula for a. Show all steps in your work.

OpenStudy (anonymous):

help please

OpenStudy (abmon98):

\[d=ut+1/2at^2 \] \[d-ut=1/2at^2 \] \[d-ut=at^2/2\] \[2d-2ut=at^2\] \[2d-2ut=at^2\] \[2d-2ut/t^2=a\]

OpenStudy (anonymous):

thank you so much @Abmon98

OpenStudy (abmon98):

your most welcome :)